# SigMA: Path Signatures and Multi-head Attention for Learning Parameters in fBm-driven SDEs

### v1

@@ 4.2.3 Ablation: convolutional and MLP layers
In this section, we investigate the impact of the convolutional and MLP layers in the SigMA model. It is noted in [tong_2023_SigFormer] that the Transformer in their model can extract representations from signatures without using a convolutional layer. Therefore, we aim to determine whether including the convolutional and MLP layers in the SigMA model improves accuracy and stability during training. We compare SigMA to three different architectures, varying the inclusion of these layers (details in Appendix A).
The results are summarized in Table 3, and a representative loss curve is shown in Figure 3. These results demonstrate that the SigMA architecture outperforms the one used in [tong_2023_SigFormer] (i.e., SigMA without convolutional layer & MLP) in terms of test error accuracy. This superiority extends across various stochastic processes and different training data settings (i.e., length and sampling of $H$). Figure 3 further suggests that SIGMA’s architecture exhibits stable training dynamics. Notably, the inclusion of the MLP appears to contribute significantly to this robustness, making the SigMA model stable and accurate. Architectures without the MLP achieve good training and test accuracy in some cases but lack consistency and can be erratic. By contrast, adding a convolutional layer appears to improve training accuracy.
@@
    SigMA without convolutional layer & MLP actually has the same architecture as the SigFormer model used in [tong_2023_SigFormer].
@@ 4.3 Hurst parameter estimation
The analysis in Section 4.2 provides a reference for the architecture of the SigMA model, which we will evaluate using single-parameter estimation problems in this section. To this end, we train a variety of neural network models, as mentioned in the beginning of Section 4, to perform the estimation using synthetic paths from Section 4.1. That is, we aim to learn the mapping $\mathbf{X}^{(n)}\mapsto H$. All the settings remain the same as described in Section 4.1.
To present the results clearly, we plot them in Figure 4, and the details can be found in Table 4. From these results, there are three conclusions to be drawn. Firstly, the SigMA model tends to achieve the best accuracy for the case of fOU and rHeston and it also achieve good accuracy, comparing with the competing models, for the case of fBm. Moreover, both the SigMA and DeepSigNet achieve high accuracy with significantly lower complexity than competing models, a performance improvement we attribute to the signature transform integrated into their architectures. In fact, the signature transform can extract core patterns from sequences, thereby filtering noise while preserving critical information for further processing. Secondly, both the SigMA and DeepSigNet models offer a significant advantage: their complexity, measured by the number of parameters, does not increase with input length. This excellent feature arises from the introduction of the signature transform, where the length of the path signature depends only on the number of features. Notably, the lifting operation before taking the signature transform would violate this advantage, therefore, as discussed in Section 4.2.1, we choose the stride equal to half the input length to avoid this issue. Thirdly, for pure fBm the test error of SigMA slightly deteriorates when increasing the number of time steps from 1000 to 1500 (Figure 4(a)), a behaviour also observed for the Transformer baseline. We attribute this to the combination of the strong self-similarity of fBm and the fixed-dimensional signature/attention bottleneck in the architecture: as the path length grows, a larger number of highly correlated increments must be compressed into a signature of fixed order, so that additional observations become largely redundant and may even degrade generalization. Similar non-monotone finite-sample effects with respect to sample size have been reported for classical Hurst estimators in long-memory settings. By contrast, the fOU and rHeston models possess additional drift and volatility structure; for these processes longer paths carry genuinely more information, and SigMA’s performance continues to improve (at least does not deteriorate) as the input length increases (Figures 4(b)-4(c)).
(a) fBm
(b) fOU
(c) rHeston
Table 4: Test RMSEs for Hurst estimation across stochastic processes of varying lengths (100100-$1500$), obtained using various NN models, are averaged over 3 training runs. Here, $H\sim\text{Uniform}(0,\,1)$ for the fBm, $H\sim\text{Uniform}(0.5,\,1)$ for the fOU, and $H\sim\text{Beta}(1,\,9)$ for the rHeston. The best accuracies across different NN models are in bold.
@@
  • 1
    The Transformer proposed by [vaswani_2017].
  • 2
    The CNN proposed by [stone2020calibrating].
  • 3
    The LSTM proposed by [csanady2024parameter].
  • 4
    The DeepSigNet proposed by [Lyons_2019_DeepSignature].
@@ 4.4.2 Numerical results
We perform two different multiple-parameter estimation experiments: the first involves simultaneous estimation of four parameters $H,\,\alpha,\,\mu$ and $\theta$ for the fOU process, and the second involves simultaneous estimation of $H,\,\kappa_{1},\,\kappa_{2}$ and $\theta$ for the rHeston process. These two experiments represent joint estimation under conditions of long-range dependence and extreme roughness, respectively. For the sake of exposition, we report the average RMSE [^9] Let $\{(x_{i},y_{i})_{i=1,\ldots,n}:x_{i},y_{i}\in\mathbb{R}^{d}\}$ be a dataset of $n$ samples, each with $d$ features. The average RMSE is then defined as $\frac{1}{d}\sqrt{\frac{1}{n}\sum\limits_{i=1}^{n}\|x_{i}-y_{i}\|_{2}^{2}}$, where $\|\cdot\|_{2}$ denotes the Euclidean norm. for all parameters combined, rather than presenting the individual RMSE values for each parameter. Furthermore, to compare the robustness of these models, we also compute the ranked average root squared error (RSE) [^10] Similarly, the average RSE of each sample is defined as $\frac{1}{d}\|x_{i}-y_{i}\|_{2}$ for $i=1,\ldots,n$. on the test dataset samples, presenting the maximum, $75\%$ quantiles and $25\%$ quantiles of the ranked average RSE for all parameters. All the results are shown in Table 6.
(a) fOU
(b) rHeston
@@ 4.5.2 Application to Li-ion battery degradation
In order to quantify the long-range dependence in the degradation dynamics, we use the Higuchi method and the rescaled range (R/S) analysis for statistical benchmarks [csanady2024parameter]. Similar to Section 4.5.1, we use the series data of the battery capacity proportion to create $46$ vectors of length $100$-cycles (entries 0-cycle to $99$-cycle, $10$-cycle to $109$-cycle and so on) to predict the Hurst parameter.
@@ 5 Conclusion
The empirical case studies further highlight the practical relevance of the proposed approach. In the calibration of equity-index realized volatility, SigMA provided accurate and stable estimates of the Hurst (Hölder) exponent with significantly lower model complexity than competing architectures. In the analysis of Li-ion battery degradation data, SigMA captured persistent long-memory effects more reliably than conventional deep learning models, producing stable estimates that align well with classical statistical benchmarks while being less sensitive to non-stationarity and short-window artifacts.

### v2

@@ 4.2.3 Ablation: convolutional and MLP layers
In this section, we investigate the impact of the convolutional and MLP layers in the SigMA model. It is noted in [31] that the Transformer in their model can extract representations from signatures without using a convolutional layer. Therefore, we aim to determine whether including the convolutional and MLP layers in the SigMA model improves accuracy and stability during training. We compare SigMA to three different architectures, varying the inclusion of these layers (details in A).
The results are summarized in Table 3, and a representative loss curve is shown in Figure 4. These results demonstrate that the SigMA architecture outperforms the one used in [31] (i.e., SigMA without convolutional layer & MLP) in terms of test error accuracy. This superiority extends across various stochastic processes and different training data settings (i.e., length and sampling of $H$). Figure 4 further suggests that SIGMA’s architecture exhibits stable training dynamics. Notably, the inclusion of the MLP appears to contribute significantly to this robustness, making the SigMA model stable and accurate. Architectures without the MLP achieve good training and test accuracy in some cases but lack consistency and can be erratic. By contrast, adding a convolutional layer appears to improve training accuracy.
We emphasize that the inclusion of the convolutional and MLP layers in SigMA is more than an architectural aggregation. They conceptually serve as an essential encoder and decoder, respectively. In particular, the convolutional layer acts as a learnable encoder that extracts local patterns and maps the raw trajectory into a latent representation. Additionally, as discussed in Remark 1, the convolutional layer also mitigates the ”curse of dimensionality”. The signature transform then provides a geometric representation of this latent path. Subsequently, the attention and MLP layers act as a decoder that maps these geometric features to parameter estimates. This ”Encoder–Geometry–Decoder” pipeline fundamentally supports the stability and accuracy observed in our experiments.
Finally, we performed a grid search over the convolutional kernel size and the hidden dimension of the MLP (see B for details) to verify that our results are robust to hyperparameter choices.
@@
    SigMA without convolutional layer & MLP actually has the same architecture as the SigFormer model used in [31].
@@ 4.3 Hurst parameter estimation
The analysis in Section 4.2 provides a reference for the architecture of the SigMA model, which we will evaluate using the Hurst parameter estimation problem in this section. To this end, we train a variety of neural network models, as mentioned in the beginning of Section 4, to perform the estimation using synthetic paths from Section 4.1. That is, we aim to learn the mapping $\mathbf{X}^{(n)}\mapsto H$. For comparison, we also introduce three classical estimators: the Detrended Fluctuation Analysis (DFA), the Wavelet estimator and the Whittle estimator. All the settings remain the same as described in Section 4.1.
We begin this section by evaluating SigMA on the Hurst parameter estimation for one-dimensional fBm-driven processes. We present the results for both NN models and classical estimators in Table 4, and plot only the NN results in Figure 5 to focus on the comparative performance of the neural architectures.
Table 4: Test RMSEs for Hurst estimation across stochastic processes of varying lengths (100100-$1500$), obtained using various NN models and classical estimators, are averaged over 3 training runs. Here, $H\sim\text{Uniform}(0,\,1)$ for the fBm, $H\sim\text{Uniform}(0.5,\,1)$ for the fOU, and $H\sim\text{Beta}(1,\,9)$ for the rHeston. The best accuracies across different NN models are in bold.
@@
DFA 100 4.28e-01 2.61e-01 4.54e-01 -
500 3.79e-01 2.03e-01 2.72e-01 -
1000 3.88e-01 1.99e-01 2.37e-01 -
1500 3.86e-01 1.93e-01 2.23e-01 -
Wavelet 100 4.15e-01 2.59e-01 2.29e-01 -
500 3.80e-01 2.15e-01 1.91e-01 -
1000 3.61e-01 1.97e-01 2.90e-01 -
1500 3.51e-01 1.92e-01 2.81e-01 -
Whittle 100 4.40e-01 2.58e-01 2.70e-01 -
500 4.06e-01 2.22e-01 1.62e-01 -
1000 4.11e-01 2.15e-01 1.38e-01 -
1500 4.06e-01 2.07e-01 1.30e-01 -
Surprisingly, the classical estimators perform poorly on our synthetic datasets compared to the NN models. However, this can be explained mathematically: the DFA requires scale windows spanning multiple decades, and the Wavelet/Whittle estimators suffer from severe variance in their lowest-frequency bins when applied to short trajectories. Moreover, the drift of the fOU and the stochastic volatility of the rHeston violate the strict global scaling and stationarity assumptions required by asymptotic statistical estimators, further limiting their performance.
For the results of NN models, there are three conclusions to be drawn. Firstly, the SigMA model tends to achieve the best accuracy for the case of fOU and rHeston and it also achieve good accuracy, comparing with the competing models, for the case of fBm. Moreover, both the SigMA and DeepSigNet achieve high accuracy with significantly lower complexity than competing models, a performance improvement we attribute to the signature transform integrated into their architectures. In fact, the signature transform can extract core patterns from sequences, thereby filtering noise while preserving critical information for further processing. Secondly, both the SigMA and DeepSigNet models offer a significant advantage: their complexity, measured by the number of parameters, does not increase with input length. This excellent feature arises from the introduction of the signature transform, where the length of the path signature depends only on the number of features. Notably, the lifting operation before taking the signature transform would violate this advantage, therefore, as discussed in Section 4.2.1, we choose the stride equal to half the input length to avoid this issue. Thirdly, for pure fBm the test error of SigMA slightly deteriorates when increasing the number of time steps from 1000 to 1500 (Figure 5(a)), a behaviour also observed for the Transformer baseline. We attribute this to the combination of the strong self-similarity of fBm and the fixed-dimensional signature/attention bottleneck in the architecture: as the path length grows, a larger number of highly correlated increments must be compressed into a signature of fixed order, so that additional observations become largely redundant and may even degrade generalization. Similar non-monotone finite-sample effects with respect to sample size have been reported for classical Hurst estimators in long-memory settings. By contrast, the fOU and rHeston models possess additional drift and volatility structure; for these processes longer paths carry genuinely more information, and SigMA’s performance continues to improve (at least does not deteriorate) as the input length increases (Figures 5(b)-5(c)).
(a) fBm
(b) fOU
(c) rHeston
As an illustration, we provide empirical measurements of computational cost for Hurst estimation of the fOU process. Figure 6 reports the measured training time for all neural architectures as a function of the input trajectory length under identical experimental conditions. The results show that SigMA exhibits substantially lower training time than transformer-based and recurrent architectures, while remaining moderately more expensive than convolutional networks due to the additional preprocessing step associated with signature computation.
For fixed truncation order, this preprocessing overhead grows approximately linearly with the sequence length. Moreover, we observe that the computational cost of SigMA remains comparable to that of other sequence models and is significantly more favorable than that of transformer architectures for the trajectory lengths considered in this study.
Figure 6: The training time (in seconds) of a typical run for Hurst estimation of the fOU process across all NN models.
Next, we test SigMA on multivariate fOU processes driven by multiple fBms [9] to explore its performance and scalability in higher dimensions. Here, we use SigMA to simultaneously estimate the Hurst parameters of multiple fBms driving the multivariate fOU processes. The results in Figure 7 verify that the input dimension has little impact on SigMA’s scalability, thanks to the convolutional preprocessing layer discussed in Remark 1. While the accuracy decreases slightly in the multivariate setting, it stays well within an acceptable limit.
(a) Parameter size vs. Dimensionality
(b) RMSE vs. Dimensionality
Figure 7: Performance of SigMA on multivariate fOU processes across different dimensions.
@@ 4.4.2 Numerical results
We perform two different multiple-parameter estimation experiments: the first involves simultaneous estimation of four parameters $H,\,\alpha,\,\mu$ and $\theta$ for the fOU process, and the second involves simultaneous estimation of $H,\,\kappa_{1},\,\kappa_{2}$ and $\theta$ for the rHeston process. These two experiments represent joint estimation under conditions of long-range dependence and extreme roughness, respectively. For the sake of exposition, we first report the average RMSE [^9] Let $\{(x_{i},y_{i})_{i=1,\ldots,n}:x_{i},y_{i}\in\mathbb{R}^{d}\}$ be a dataset of $n$ samples, each with $d$ features. The average RMSE is then defined as $\frac{1}{d}\sqrt{\frac{1}{n}\sum\limits_{i=1}^{n}\|x_{i}-y_{i}\|_{2}^{2}}$, where $\|\cdot\|_{2}$ denotes the Euclidean norm. for all parameters combined. Furthermore, to compare the robustness of these models, we also compute the ranked average root squared error (RSE) [^10] Similarly, the average RSE of each sample is defined as $\frac{1}{d}\|x_{i}-y_{i}\|_{2}$ for $i=1,\ldots,n$. on the test dataset samples, presenting the maximum, $75\%$ quantiles and $25\%$ quantiles of the ranked average RSE for all parameters. All the results are shown in Table 6.
@@
(a) fOU
(b) rHeston
For further insight into the joint estimation results, Table 10 provides the RMSE and variance for each individual parameter. These results reveal that parameters governing short-term fluctuations (such as volatility-related parameters and the Hurst exponent) tend to be estimated more accurately than drift or mean-reversion parameters. A possible explanation is that volatility and roughness properties primarily influence local increments of the trajectory, which are effectively captured by low-order signature features. In contrast, drift and mean-reversion parameters influence the longer-term evolution of the process and therefore require longer observation horizons for accurate estimation [23, 17, 14].
@@ 4.5.2 Application to Li-ion battery degradation
In order to quantify the long-range dependence in the degradation dynamics, we use the Higuchi method and the rescaled range (R/S) analysis for statistical benchmarks [8]. Similar to Section 4.5.1, we use the series data of the battery capacity proportion to create $46$ vectors of length $100$-cycles (entries 0-cycle to $99$-cycle, $10$-cycle to $109$-cycle and so on) to predict the Hurst parameter.
@@
The degradation series exhibits strong non-stationarity, including long-term trends and local regime shifts. Classical estimators such as R/S analysis and Higuchi’s method rely on global scaling assumptions and are therefore sensitive to these effects. For example, the R/S estimator is known to overestimate the Hurst exponent in the presence of deterministic trends. Signature-based models operate differently because they rely on geometric representations of path increments. The signature transform captures the ordered structure of increments rather than absolute values of the signal. In addition, the convolutional preprocessing layer extracts local patterns from the raw trajectory before the signature transform is applied. This combination makes the model less sensitive to isolated spikes or local irregularities in the observations.
@@
4.6 Summary of architectural comparisons
Based on the results presented in Sections 4.2-4.5, we summarize the differences between SigMA and comparable models in terms of architectural designs, applicable scenarios, and performance.
SigFormer applies attention mechanisms directly to signature features and uses a linear output layer. In contrast, SigMA introduces a convolutional preprocessing layer before the signature transform and an MLP after the attention layer. Empirically, we observe that replacing the final linear layer with an MLP improves both training stability and predictive accuracy (see Figure 4), suggesting that the mapping from geometric signature features to model parameters is highly nonlinear.
DeepSigNet directly feeds signature features into an MLP, whereas SigMA introduces a multi-head self-attention layer between the signature transform and the MLP. In our architecture, different attention heads process signature features associated with different truncation levels, allowing the model to weight geometric information arising from different interaction orders. In the joint multi-parameter estimation (Section 4.4), DeepSigNet tends to over-compress the geometric information, leading to more outliers and higher quantiles (see Table 6). In contrast, by weighting the importance of different signature levels, SigMA can more effectively capture different geometric features of paths, leading to lower estimation errors and higher stability.
Finally, classical models such as CNN, LSTM, and Transformer operate directly on raw paths instead of path signatures. For long sequences, these models need significantly more parameters to capture temporal dependencies, making them expensive to train. As shown in Table 4, for an input length of $1500$ steps, the LSTM and Transformer models require over 2424M and 7.57.5M parameters, respectively. In contrast, SigMA keeps a small, constant parameter count ($87,226$) regardless of the sequence length, achieving a better balance between model complexity and accuracy.
@@ 5 Conclusion
The empirical case studies further highlight the practical relevance of the proposed approach, particularly its ability to overcome the distribution mismatch between simulated training data and real-world observations. In the calibration of equity-index realized volatility, SigMA provided accurate and stable estimates of the Hurst (Hölder) exponent with significantly lower model complexity than competing architectures. In the analysis of Li-ion battery degradation data, SigMA captured persistent long-memory effects more reliably than conventional deep learning models, producing stable estimates that align well with classical statistical benchmarks while being less sensitive to non-stationarity and short-window artifacts. This generalizability likely stems from signature representations capturing the structural properties of trajectories rather than specific marginal distributions.
@@
While SigMA generally works well, it struggles at theoretical boundaries of the Hurst parameter. As $H\to 0$, paths become extremely rough, and the key details shift to higher-order signature terms. By defaulting to $N=3$, SigMA loses this information and reduces in accuracy, while increasing $N$ is too computationally expensive. At the other extreme ($H\approx 0.5$), the process is memoryless. SigMA’s self-attention mechanism isn’t needed here and may actually overfit by searching for nonexistent long-term dependencies. For such cases, simpler classical estimators are a better choice.
@@ Appendix A Architectural details of baseline and signature-based neural networks
Appendix B Hyperparameter grid search for convolutional and MLP layers
Table 9: Test RMSEs for Hurst estimation across stochastic processes with length 100, obtained using SigMA under different combinations of convolutional kernel sizes and MLP hidden layer sizes. Results are averaged over 3 training runs. Here, $H\sim\text{Uniform}(0,\,1)$ for the fBm, $H\sim\text{Uniform}(0.5,\,1)$ for the fOU, and $H\sim\text{Beta}(1,\,9)$ for the rHeston. The default configuration is marked.
Convolutional kernel size MLP hidden layer size Stochastic processes #\#Params
fBm fOU rHeston
2 16 1.68e-02 3.29e-02 5.58e-02 79095
32 1.69e-02 3.14e-02 5.40e-02 87223
64 1.76e-02 2.93e-02 5.32e-02 109623
3 (Default) 16 1.57e-02 2.82e-02 5.51e-02 79098
32 (Default) 1.49e-02 2.70e-02 5.01e-02 87226
64 1.38e-02 2.55e-02 4.56e-02 109626
4 16 1.55e-02 2.64e-02 5.64e-02 79101
32 1.47e-02 2.72e-02 5.30e-02 87229
64 1.60e-02 3.09e-02 3.99e-02 109629
Appendix C Individual parameter results for joint multi-parameter estimation
Table 10: Test results for multiple parameters estimation across stochastic processes with length 500, obtained using various NN models, are based on $10$ training runs. Here, numbers in parentheses represent the variance of the RSEs. The best accuracies across different NN models are in bold.
Models RMSEs for fOU RMSEs for rHeston #\#Params
$H$ $\alpha$ $\mu$ $\sigma$ $H$ $\kappa_{1}$ $\kappa_{2}$ $\theta$
Transformer 0.165 0.764 0.149 0.664 0.086 0.993 0.130 0.416 2558005
(1.12e-2) (2.96e-1) (8.28e-3) (1.49e-1) (3.27e-3) (3.71e-1) (4.86e-3) (8.71e-2)
SigMA 0.120 0.761 0.139 0.643 0.071 0.970 0.110 0.412 87341
(5.38e-3) (2.51e-1) (5.75e-3) (1.56e-1) (3.75e-3) (3.38e-1) (4.03e-3) (8.40e-2)
DeepSigNet 0.451 1.228 0.741 0.998 0.084 0.992 0.100 0.414 9360
(2.25e-1) (9.92e-1) (7.50e-1) (4.63e-1) (3.16e-3) (3.57e-1) (3.67e-3) (8.09e-2)
CNN 0.129 1.051 0.128 0.648 0.078 1.241 0.088 0.587 501220
(5.13e-3) (3.28e-1) (4.60e-3) (1.64e-1) (2.69e-3) (5.62e-1) (2.56e-3) (1.29e-1)
LSTM 0.158 0.977 0.212 1.030 0.107 1.275 0.157 0.909 8383299
(9.33e-3) (4.43e-1) (1.70e-2) (3.70e-1) (4.67e-3) (5.74e-1) (9.51e-3) (2.82e-1)